# Federated Learning in Adversarial Environments: Testbed Design and Poisoning Resilience in Cybersecurity

### v1

@@ III-B Poisoned Model Performance Results
Figure 3 presents the performance of each client across three experiments with different numbers of clients (3, 4, and 5) in federated learning settings. The evaluation metrics displayed include evaluation accuracy, evaluation loss, and F1-score over 20 rounds of training. Each experiment involves a unique configuration of clients, with client 3 being consistently poisoned across all trials. The figures provide valuable insights into how the introduction of poisoned data from client 3 affects the global model’s overall performance and how the system’s robustness changes as the number of clients increases.
@@
As the number of clients increases to 4 and 5, the overall evaluation loss and performance trends suggest that federated learning becomes more robust to individual client poisoning. Clients 1, 2, 4, and 5 show improvements in both accuracy and F1-score, even with the presence of a poisoned client. This resilience can be attributed to the aggregation of updates from a larger number of benign clients, which helps to dilute the effect of client 3’s poisoned updates. The results suggest that while poisoned clients can have a severe impact on local model performance, the collective aggregation of updates in federated learning helps maintain global model accuracy, especially as the number of clients increases. However, it also emphasizes the importance of incorporating defense mechanisms to detect and mitigate poisoned contributions, particularly in smaller federated learning environments where the influence of a poisoned client is more pronounced.
In Figure 3, we observe that Client 3 consistently exhibits the worst performance across all three metrics -evaluation accuracy, loss, and F1-Score- in all three experiments conducted. This trend is expected, as approximately 55% of Client 3’s data was heavily poisoned. The graphs have highlighted the impact of poison data on client 3’s performance.
@@ III-C Comparison of Aggregated Models
In Figure 4, the performance of the federated learning system is compared under two conditions: a normal scenario (top row) and a poisoned scenario (bottom row). The results are presented across three metrics: evaluation accuracy, evaluation loss, and F1-score, for configurations of 3, 4, and 5 clients. The overall trends in the graphs show how the model performs during the 20 rounds of training under both normal and poisoned conditions.
In the normal scenario (top row), the evaluation accuracy shows a steady increase during the first 5 rounds, stabilizing around 0.7 for all client configurations (3, 4, and 5 clients). The evaluation loss also significantly decreases over time, with a sharp drop in the first 5 rounds before leveling out as the model converges. The F1-score follows a similar pattern, increasing rapidly in the first few rounds and stabilizing at around 0.7. The similar behavior of the different client configurations in terms of accuracy, loss, and F1-score indicates that in the absence of poisoning, the federated learning model performs consistently well, regardless of the number of clients involved.
In contrast, the poisoned scenario (bottom row) exhibits a notable degradation in performance, especially in evaluation accuracy and F1-score. While the initial increase in accuracy during the first few rounds appears similar to the normal scenario, the final accuracy levels off at a slightly lower value compared to the normal condition. Evaluation loss also shows a similar decreasing trend, but it does not reach as low a value as in the normal condition, indicating that the model struggles to fully converge when poisoned data is introduced. Additionally, the F1-score, while following the same upward trajectory, stabilizes at a marginally lower value compared to the normal scenario. This suggests that while the model remains relatively robust, the presence of poisoned data still impacts its overall classification performance.
These results demonstrate that while federated learning systems can maintain some level of resilience to poisoned data, the performance in terms of accuracy and F1-score is slightly reduced. The model does not fully converge as effectively as in the normal scenario, highlighting the need for further improvements in defense mechanisms to mitigate the impact of adversarial attacks like data poisoning.
In Figure 4, despite the significant performance differences between the poisoned client and the other clients, the global model’s overall performance remains largely unaffected.
@@ IV Conclusion and Future Work
In this paper, we developed and evaluated a federated learning testbed to assess the resilience of models in the face of adversarial data poisoning, specifically in a cybersecurity use case involving DNP3 intrusion detection. Through our experiments with both normal and poisoned data, we demonstrated that while federated learning systems show robust performance under normal conditions, poisoned data significantly impacts local model accuracy and convergence. However, as the number of clients increases, the global model remains more resilient due to the aggregation of updates from non-poisoned clients. The use of the Flower framework proved highly effective for its flexibility, scalability, and compatibility with resource-constrained devices like the Nvidia Jetson Nano. Our findings underscore the need for robust defense mechanisms to address federated learning’s susceptibility to adversarial attacks in critical applications like cybersecurity. For future work, it is possible to expand the testbed to evaluate defense strategies such as Byzantine Robust Aggregation and differential privacy, aiming to mitigate the impact of poisoning and enhance system robustness. Additionally, we plan to explore more complex attack strategies to further stress-test federated learning models and inform the development of secure and resilient FL systems for privacy-sensitive environments.

### v2

@@ III-B Poisoned Model Performance Results
Figure 3 presents the performance of each client across three experiments with different numbers of clients (3, 4, and 5) in FL settings. The evaluation metrics displayed include evaluation accuracy, evaluation loss, and F1-score over 20 rounds of training. Each experiment involves a unique configuration of clients, with client 3 being consistently poisoned across all trials.
@@
As the number of clients increases to 4 and 5, the overall evaluation loss and performance trends suggest that FL becomes more robust to individual client poisoning. Clients 1, 2, 4, and 5 show improvements in both accuracy and F1-score, even with the presence of a poisoned client. This resilience can be attributed to the aggregation of updates from a larger number of benign clients, which helps to dilute the effect of client 3’s poisoned updates. The results suggest that while poisoned clients can have a severe impact on local model performance, the collective aggregation of updates in FL helps maintain global model accuracy, especially as the number of clients increases. However, it also emphasizes the importance of incorporating defense mechanisms to detect and mitigate poisoned contributions, particularly in smaller FL environments where the influence of a poisoned client is more pronounced.
@@ III-C Comparison of Aggregated Models
In Figure 4, the performance of the FL system is compared under two conditions: a normal scenario (top row) and a poisoned scenario (bottom row). The results are presented across three metrics: evaluation accuracy, evaluation loss, and F1-score, for configurations of 3, 4, and 5 clients.
In the normal scenario (top row), the evaluation accuracy shows a steady increase during the first 5 rounds, stabilizing around 0.7 for all client configurations (3, 4, and 5 clients). The evaluation loss also significantly decreases over time, with a sharp drop in the first 5 rounds before leveling out as the model converges. The F1-score follows a similar pattern, increasing rapidly in the first few rounds and stabilizing at around 0.7. The similar behavior of the different client configurations in terms of accuracy, loss, and F1-score indicates that in the absence of poisoning, the FL model performs consistently well, regardless of the number of clients involved.
In contrast, the poisoned scenario (bottom row) exhibits a notable degradation in performance, especially in evaluation accuracy and F1-score. While the initial increase in accuracy during the first few rounds appears similar to the normal scenario, the final accuracy levels off at a slightly lower value compared to the normal condition. Evaluation loss also shows a similar decreasing trend, but it does not reach as low a value as in the normal condition, indicating that the model struggles to fully converge when poisoned data is introduced. Additionally, the F1-score, while following the same upward trajectory, stabilizes at a marginally lower value compared to the normal scenario.
These results show that while FL systems can maintain some level of resilience to poisoned data, the performance in terms of accuracy and F1-score is slightly reduced. The model does not fully converge as effectively as in the normal scenario, highlighting the need for further improvements in defense mechanisms to mitigate the impact of adversarial attacks like data poisoning.
@@ IV Conclusion and Future Work
In this paper, we developed and evaluated a FL testbed to assess the resilience of models in the face of adversarial data poisoning, specifically in a cybersecurity use case involving DNP3 intrusion detection. Through our experiments with both normal and poisoned data, we demonstrated that while FL systems show robust performance under normal conditions, poisoned data significantly impacts local model accuracy and convergence. However, as the number of clients increases, the global model remains more resilient due to the aggregation of updates from non-poisoned clients. The use of the Flower framework proved highly effective for its flexibility, scalability, and compatibility with resource-constrained devices like the Nvidia Jetson Nano. Our findings underscore the need for robust defense mechanisms to address FL’s susceptibility to adversarial attacks in critical applications like cybersecurity.
Future work includes implementing and evaluating specific defenses, such as various Byzantine Robust Aggregation algorithms (e.g., distance-based Krum, statistics-based Median) discussed in [13]. We also plan to integrate differential privacy by adding calibrated noise to client updates before aggregation, assessing the privacy-utility trade-off as analyzed in other frameworks [14]. Additionally, we plan to explore more complex attack strategies to further stress-test FL models and inform the development of secure and resilient FL systems for privacy-sensitive environments.